At Lequals66 , Kequals130 , the marginal product of labor is 4 and the marginal product of capital is 2. What is the marginal rate of technical substitution (labor measured on the horizontal axis )? The marginal rate of technical substitution is nothing. (Enter a numeric response using a real number rounded up to two decimal places.)
Answer #1

Marginal rate of technical substitution=MPL/MPK

MRTS=4/2=2

MRtS=2, means 2 unit of capital can be sacrificed in order to get an additional unit of labor keeping total output same